### v1

# Strain-Enhanced Altermagnetism in Ca3Ru2O7

Abstract
Perovskite compounds exhibit enhanced octahedral distortions coupled with strong electronic correlations, providing a promising platform to explore and tune altermagnetic (AM) order. In this work, we investigate AM phases in Ca3Ru2O7, a well-known perovskite that hosts antiferromagnetism coupled to structural degrees of freedom. We demonstrate that Ca3Ru2O7 is a Kramers antiferromagnet in its ground state. However, a Néel-type magnetic configuration reveals a P-2 d-wave AM, hosting orbital-selective altermagnetism analogous to Ca2RuO4. We further explore the effects of biaxial strain on the stability between the antiferromagnetic ground state and the AM phase. Our results suggest that the AM phase becomes more stable than the AFM phase below -2% compressive strain. Additionally, strain can either enhance or suppress AM bands, with enhancements reaching up to 9% under tensile strain. To quantify the AM band splitting, we introduce the ”Altermagnetic Merit Figure” and analyze the role of electronic localization, delocalization, and octahedral distortions in AM behavior and magnetic stability changes under strain.
1 Introduction
[…]
Due to the breaking of time-reversal symmetry, the spin-orbit interaction can induce the so-called weak ferromagnetism [6] in altermagnet via a canting of the spins or other mechanisms depending on the crystal symmetries of the material [7, 8, 9, 10, 11, 12].
Given these findings, it is natural to explore whether strain could also play a key role in tuning AM order in other correlated materials. In this regard, Ruddlesden-Popper perovskites, which host diverse quantum states, could represent an interesting class of systems for investigating strain-induced AM behavior. Since many of their electronic phases are closely linked to structural distortions, these perovskites offer a promising platform to study the emergence of AM phases in relation to their inherent structural and electronic degrees of freedom [13, 14, 15, 16].
Ruddlesden-Popper compounds of the form (Ca,Sr)n+1RunO3n+1 are well-known correlated systems whose intriguing electronic properties are strongly influenced by their symmetries and octahedral distortions. These distortions, including rotations, Jahn-Teller effects, and polar distortions, can break symmetries that transform opposite spin sublattices, potentially enabling the emergence of AM states [14]. Over the past decade, these materials have attracted significant attention due to their exotic electronic phases, including Mott insulating behavior [17], metamagnetism [18], and unconventional superconductivity [19]. Revisiting their magnetic phases in the context of AM order could provide new insights into phase transitions between ferromagnetism, antiferromagnetism, and AM, offering a framework to understand the interplay between symmetry, octahedral distortions, and magnetism in correlated systems.
Among the (Ca,Sr)n+1RunO3n+1 series, we focus on Ca3Ru2O7 (CRO) to explore AM states. CRO is a fascinating compound that undergoes magnetic AFM transitions with temperature [20, 21, 22]. Over the past decade, experiments have shown that its electronic phases can be manipulated by magnetic fields [23], pressure [24], and, more recently, strain [25]. This raises the question of whether CRO exhibits altermagnetism or under what conditions its AFM properties can be tuned toward an AM state. Moreover, very few altermagnetic materials lack inversion symmetry[26, 27]. Tuning altermagnetism in Ca3Ru2O7 would introduce another noncentrosymmetric AM material, enabling the study of the interplay between non-relativistic and relativistic spin-splitting.
In this work, we explore various magnetic configurations in the CRO system to identify those that exhibit AM order. We begin by analyzing the electronic properties and then discuss the symmetry characteristics that give rise to AM phases in CRO. Furthermore, we examine the system under strain to determine the conditions under which the AM phase can be stabilized. Our results demonstrate that the altermagnetic character can be enhanced or suppressed under biaxial tensile and compressive strain. We discuss the role of octahedral distortions and the localization/delocalization effects induced by the strain in the changes to altermagnetic band splitting. Additionally, we introduce the altermagnetic quantity (AMQ) as the “figure of merit” to quantify the degree of enhancement or suppression of the band splitting with respect to the non-strained system.
2 Computational details
We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling (SOC). We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[28] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values. The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334. We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation. We use a fine k-grid of 14147 within the tetrahedron method for the density of states. We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
To account for the electronic correlation effects in the -orbitals of Ru atoms, we apply a Hubbard on-site Coulomb parameter using the Dudarev approximation [29] with U = 1.0 eV, a sufficiently accurate value for computing the electronic and structural properties of CRO, as reported in our previous study [30].
3 Results
3.1 Exploring AM states
CRO exhibits a non-centrosymmetric crystal structure with the space group Bb21m [21]. The associated point group is mm2, which includes a twofold rotation axis along the -axis (C2z), two perpendicular mirror planes parallel to the and planes, and a helicoidal translation along (E, C2z, , , 21). This point group has been associated with the exhibition of altermagnetic (AM) phases [31]. The magnetic moment arises predominantly from the Ru 4 electrons, in which, at low temperatures, the Ru atoms order ferromagnetically within the bilayers while aligning antiferromagnetically between them, forming an A-type antiferromagnetic structure where the spins are oriented along the -axis (see Fig. 1). Given this A-type antiferromagnetic order, the primitive cell of CRO contains spin-up sites, with a translational symmetry operation that maps spin-up into spin-down sites (along the -axis). Consequently, the ground state is not altermagnetic (see Fig. 1) [32].
To explore possible AM states and their stability, we analyzed CRO under different magnetic configurations. Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [34], labeled as configurations A, B, C, and D (see Fig.2). Table 1 shows the energy differences relative to the ground-state configuration A, as well as some relevant structural, electronic, and magnetic properties for our study. Our results indicate that AM states appear in configurations B and C, where configuration B has an antiferromagnetic spin arrangement along the - and -axes, while configuration C shows this along the -axis. Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [35], and in the sister compound Sr3Ru2O7 (configuration D) [36].
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
E | E-State | M-Phase | () | |
A | 0 | Metallic | AFM | 1.44 |
B | 34.90 | N-Insulator | AM | 1.35 |
C | 36.46 | Metallic | AM | 1.38/1.45 |
D | 39.22 | N-Insulator | AFM | 1.35 |
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
In configuration B, the magnetic moments adopt a Néel-type ordering, where spin-up and spin-down orientations alternate within the -plane and along the -axis, forming spin-singlet wave functions.
This configuration promotes spatial delocalization, which is characteristic of bonding-like states. Since bonding states enhance wave function overlap, that is, increase the spatial proximity of electrons, configuration B is more sensitive to the Hubbard- term.
As a result, a band gap opens at relatively low U values, specifically . In contrast, configuration A remains metallic under the same conditions. Further analysis is provided in the Supplemental Material (Section SII).
These contrasting results demonstrate CRO’s highly correlated electronic nature, driven by the tight interplay between magnetic and structural degrees of freedom [30, 37, 38, 39]. Finally, we confirm that these characteristics are predominantly electronic, arising from the magnetic spin configuration, and are independent of volume changes or structural distortions.
We now focus on the electronic band structure of both systems. Figure 3 shows the total and partial density of states for configurations A and B. Configuration A exhibits metallic behavior, while configuration B displays a narrow gap at the Fermi level. The Ru ions, located in an octahedral environment formed by oxygen atoms, are subjected to a crystal field that splits the Ru 4 levels, raising the eg states and lowering the t2g manifold energies. In configuration B, the orbitals are nearly filled, leading to a small gap in the half-filled and bands. Notably, the orbitals are the most affected, as the spin rearrangement occurs within the -plane. This behavior is similar to that observed in Ca2RuO4
[40].
Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations. The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect spin-orbit coupling (SOC) effects [41, 30]. In contrast, configuration B displays a semimetallic character and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig.4(b)). The shaded regions on the band structure indicate regions in -space where the altermagnetic spin splitting is maximal. The detailed BZ for these paths is shown on the right Fig. 4.
The presence of the non-relativistic spin-splitting is independent of the value of ky. According to the notation on the spin-momentum locking, this is P-2 d-wave altermagnets (where the d-wave is in the k-space) [32], since it is independent on ky this is a P-2 dxz-wave. These paths, which extend beyond the plane, have not been explored in previous studies, where ARPES experiments primarily focused on the plane [41]. In the next section, we discuss the preference for a specific direction in the AM character.
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, showing negligible spin splitting (see, Fig.LABEL:fig:S2). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [40]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.LABEL:fig:S2.
CRO, due to its non-centrosymmetric structure, is expected to exhibit pronounced SOC effects, as demonstrated in previous studies on configuration A (the experimental magnetic phase), where the spins are oriented along the -axis [41, 30]. In configuration B, the anisotropy also lies primarily along the -axis, with a minor component along the -axis, specifically . Figure LABEL:fig:S3 in the SM shows the band structure with SOC, by spin-component contributions and orbitals. Since the magnetic moment aligns along the -axis, a strong Kramers spin degeneracy is evident along this direction. However, spin polarization components are also observed along the -axis. The main contributors to the AM character are the dxy orbitals (from -0.7 to -0.5 eV) and the dyz orbitals (from -1.5 to -1.0 eV) (see SM, Fig. LABEL:fig:S2). The slight enhancement of the dxy contribution arises from the band splitting in the band structure induced by SOC, on top of the non-relativistic spin splitting. This induces both splitting and mixing of the spin character of the bands.
3.2 Symmetry analysis
To understand the role of symmetry in AM-CRO, consider applying a time-reversal operator (TR), which in practical terms consists of an exchange between spin-up and spin-down (see Fig. 5(a)-(b). Due to the distinct symmetry properties of many altermagnets, the original magnetic configuration cannot be recovered using only rotational symmetry operations. Following a time-reversal operation, a half-unit cell translation along the -axis, shifting (Fig. 5(c)), is combined with a screw symmetry operation by inverting and adding a half-unit cell translation along the -axis (Fig. 5(d)-(e)).
In summary, the operation z becomes necessary (Fig. 5(f)). These combined nonsymmorphic operations result in the -dependent splitting states [32]. Due to the symmetry operations that connect the spins, which are determined by rotations and translations along the and directions while leaving the direction unchanged, the AM character is found along these axes, as shown in Fig. 4. In the case of the C configuration, the symmetry operation ( +, + , - + ) becomes necessary to achieve AM order.
3.3 Tuning AFM to AM phases
To analyze the stability of configurations A and B, we calculated their total energies under various biaxial strain conditions (Fig. 6(a)). Configuration B shows the lowest energy within a compressive strain below . Fig. 6(b) shows the -lattice parameter for each strain. These results suggest a potential magnetic transition from AFM to AM states in CRO.
Next, we focus on investigating the effects of compressive (-) and tensile (+) strain on the AM features of CRO. Fig. 7 shows the band structure of configuration B under strains of = -2%, 0, and 2% (biaxial). It can be observed that compressive strain suppresses the AM features among the energy range [-2:-1] eV (compare Fig. 7 (a) with Fig.7(b)). Moreover, the tensile strain enhances the splitting being shifted to higher energies.
To quantify the effect of strain on the altermagnetic features of CRO, we calculate the average difference between spin-up and spin-down eigenvalues for each occupied band along the entire -path. Additionally, we define the altermagnetic quantity (AMQ) as the altermagnetic figure of merit, which integrates over the Brillouin zone. The AMQ is a numerical measure of the overall altermagnetic character of a system, providing a quantitative basis for comparing different strain conditions.
In our implementation, we read the spin-resolved eigenvalues from the EIGENVAL file generated by VASP, then compute the -dependent average splitting as the cumulative sum over occupied states along a particular set of k points (over a -path or -grid), given by:
where Nk is the numbers of bands at -point within specified energy range and Nocc is the occupied bands.
Note that the integration to obtain the AMQ can be performed either over the entire Brillouin zone or along a specific -path. We have verified that both methods yield similar solutions up to a constant multiplicative factor, which cancels out when percentage changes are considered.
We sum up all the occupied states rather than comparing only a few selected bands. This approach proves more reliable because the most strongly split bands can vary across different systems and even within the same material under strain, depending on details such as band character and crystal symmetry. Consequently, integrating over the entire set of occupied bands furnishes a more robust assessment of the altermagnetic spin-splitting.
The AMQ quantifies the spin-splitting across momentum space by integrating :
in our code, Simpson’s rule is used for numerical integration. When points are expressed in inverse angstroms (), the resulting AMQ is in units of . Furthermore, this straightforward procedure for quantifying the altermagnetic character may fail to correctly identify the spin-splitting in cases where bands cross or overlap. A potentially improved approach would incorporate a disentanglement method based on band character analysis, for instance, using Wannier functions or explicitly tracking the orbital character of each band. Such an approach would enable a more rigorous and accurate splitting assignment; however, its detailed implementation is beyond the scope of the present study and remains a subject for future work.
Since our primary goal is to evaluate strain effects on the altermagnetic state, we compare the AMQ under strain to the unstrained scenario. We define the percent variation as:
This relative measure properly captures how strain modifies the degree of altermagnetic spin-splitting throughout the Brillouin zone.
Fig. S9 illustrates the color-map along the entire -path (see the BZ in Fig.4) and by occupied bands (number labels at the -axis) corresponding to each strain value. The splitting by bands clearly varies as a function of the applied strain, with some bands becoming more or less altermagnetic. Under tensile strain (up to 4% of the maximum value explored), increases significantly compared to the unstrained system. By integrating using Eq. 2, we obtain AQM values of 0.0010, 0.0012, and 0.0013 for strains of -2%, 0%, and 2%, respectively. Fig.6(c) shows the percentual variation of AQM across different strain values given by Eq. 3. Under biaxial strain, the band splitting increases by approximately under tensile strain (at 2%) and decreases by about under compressive strain (at -2%).
The non-relativistic spin-splitting in Ca3Ru2O7 arises from the interplay between hopping parameters and the on-site energy difference between majority and minority -electrons. Strain introduces competing effects that influence these parameters: compression along one axis promotes electronic delocalization, weakening magnetism and reducing the AM band splitting, whereas tensile strain favors localization, enhancing magnetic exchange and increasing the AM band splitting.
From the energy–strain relationship shown in Fig. 6(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions so that configuration B is stabilized relative to configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (see, Fig.S8 SM). This effect is reflected in the rising AMQ values in Fig. 6(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.
To further assess the role of structural distortions, we performed an additional set of calculations where the atomic positions remained fixed at their zero-strain configuration. At the same time, only the lattice parameters were modified (fixed structure (FS) calculations). The results, depicted by the FS line in Fig. 6(c), reveal a similar overall trend: AMQ increases under tensile strain and decreases under compressive strain. However, for , the deviation between the relaxed and non-relaxed calculations becomes more pronounced, suggesting that octahedral distortions () play a secondary but non-negligible role in the strain response of the AM phase. Specifically, as seen in Fig. 6(d), increases under compressive strain for configuration B, further enhancing localization effects that help sustain the AM band splitting.
A direct comparison between relaxed structure and the non-relaxed structure allows us to disentangle the effects of strain-induced delocalization/localization from those arising purely due to octahedral distortions. In particular, the density of states (DOS) in Fig. S8 confirms that compressive strain broader states, reducing AM features. In contrast, tensile strain narrows the bandwidth, reinforcing the AM character. Moreover, the band structures shown in Fig. S10 illustrate how freezing the atomic positions primarily alters the gap structure rather than the AM spin splitting itself, reinforcing the observation that the observed strain-dependent AM behavior is primarily dictated by electronic exchange interactions, with octahedral distortions acting as a secondary fine-tuning mechanism.
Furthermore, an interesting behavior can be observed in the evolution of the distortion angle () as a function of strain (see Fig. 6(d)). In configuration A, increases monotonically from to strain. However, in configuration B, exhibits a more complex response, increasing under compressive strain () but displaying a non-monotonic behavior under tensile strain. This difference originates from the distinct in-plane magnetic configurations of the two phases. Compressive strain generally enhances electronic delocalization, which tends to weaken magnetic exchange interactions and favor a more symmetric crystal structure. In the case of conf. A, which is metallic (see, Fig. 3a)), the effect of in-plane strain gradually increases delocalization, reducing structural distortions and stabilizing a more symmetric octahedral environment. Conversely, electronic states are more localized in configuration B (see, Fig. 3b)), where the in-plane magnetic alignment is AFM. As a result, under compression, the system seeks to accommodate strain by increasing octahedral distortions (), effectively enhancing localization to maintain magnetic interactions, resulting in a more distorted structure than configuration A. Under tensile strain, however, competing effects emerge: while in-plane expansion reduces the need for further distortions, the slight contraction along the -axis counterbalances this effect, leading to either an increase or decrease in , depending on the precise strain conditions.
The strain-dependent behavior of is closely tied to the observed changes in the energy profile in Fig. 6(a), where configuration B becomes energetically favorable at compressive strain values above . The stabilization of the AM phase coincides with an increase in , suggesting that octahedral distortions play a crucial role in the strain-driven AFM-to-AM transition.
4 Final Remarks
In this work, we explore the emergence of altermagnetic states in the correlated material Ca3Ru2O7. While its ground state lacks AM features, alternative magnetic configurations enable their realization. AM arises in spin arrangements connected through rotational and translational symmetries. In particular, configuration B, which alternates spin orientation both in-plane and between layers, exhibits spin connectivity via rotations along the -axis and translations in the - plane. This configuration hosts a P-2 -wave altermagnetic state, which is orbital-selective, with AM primarily arising from the - orbitals. When the spin-orbit coupling is included, with the Néel vector oriented along the -axis, time-reversal symmetry breaking leads to weak ferromagnetism along the -axis due to the staggered Dzyaloshinskii-Moriya interaction [7, 15]. While configuration B has not been experimentally observed in pristine Ca3Ru2O7, it has been reported in doped systems, such as Ca3(TixRu1-x)2O7, suggesting that this system could host an AM ground state under suitable conditions. Additionally, our results indicate that strain plays a crucial role in tuning AM properties. Tensile strain along the -axis enhances AM band splitting by approximately (for strain), whereas compressive strain suppresses it by . Moreover, we predict a strain-driven phase transition from an AFM to an AM state (configuration A to B) under compressive strain exceeding . While previous experiments have manipulated magnetic anisotropy using uniaxial strain (), the impact of biaxial strain remains largely unexplored.
Strain-tuned AM phases could significantly expand the known quantum phenomena in the Can+1RunO3n+1 family, which already exhibit strain-driven transitions, such as Mott phase suppression in Ca2RuO4 [17], the emergence of a Kondo effect, and a metal-to-semiconductor transition in CaRuO3 ( strain) [42]. Our findings provide a fundamental step toward uncovering possible AM states in the Ca3Ru2O7 family and pave the way for future theoretical and experimental investigations of biaxial strain as a tool for engineering AM phases in correlated oxides.
Acknowledgments
A.L. thanks the Technische Universität Dresden Professor Fellowship Program. C. A. acknowledges the useful discussion with G. Cuono and A. Fakhredine about altermagnetism in Ca2RuO4. C. A. is supported by the Foundation for Polish Science project “MagTop” no. FENG.02.01-IP.05-0028/23 co-financed by the European Union from the funds of Priority 2 of the European Funds for a Smart Economy Program 2021–2027 (FENG). J.W.G. acknowledges financial support from Chilean ANID-FONDECY grants N. 1220700 and 1221301. T.B. would like to thank the German Science Foundation (DFG) for supporting this work via the German-Israeli Project Cooperation with the Grant No. HE 3543/42-1.